# Comparing Task Graph Scheduling Algorithms: An Adversarial Approach

### v2

@@ 6.1. Experimental Setup
One of the largest communities of researchers that depend on efficient task scheduling algorithms is that of scientists that use scientific workflows for simulation, experimentation, and more. Scientific workflows are typically big-data scale task graphs that are scheduled to run on cloud or super computing platforms. These researchers typically have little to no control over how their workflows are scheduled, instead relying on Workflow Management Systems (WFMS) like Pegasus (Deelman et al., 2015), Makeflow (Albrecht et al., 2012), and Nextflow (Di Tommaso et al., 2017) (to name a just few examples) to handle the technical execution details. For this reason, it is especially important for WFMS developers/maintainers (who choose which scheduling algorithms their system uses) to understand the performance boundaries between different algorithms for the different types of scientific workflows and computing systems their clients use.
@@
SAGA currently has datasets based on nine real-world scientific workflows (blast, bwa, cycles, epigenomics, 1000genome, montage, seismology, soykb, and srasearch). These applications come from a very wide range of scientific domains — from astronomy (montage builds mosaics of astronomical imagery) to biology (1000genome performs human genome reconstruction) to agriculture (cycles is a multi-crop, multi-year agroecosystem model). For each workflow, the runtime of each task, input/output sizes in bytes, and speedup factor (compute speed) for each machine are available from public execution trace information [^2] https://github.com/wfcommons/pegasus-instances, https://github.com/wfcommons/makeflow-instances. The inter-node communication rate, however, is not available. We set communication rates to be homogeneous so that the average CCR, or average data sizecommmunication $\frac{\text{average data size}}{\text{commmunication strength}}$, is $\frac{1}{5},\frac{1}{2},1,2,$ or $5$ (resulting in five different experiments for each workflow).
@@
Because these scientific workflows are much larger than the workflows used in the experiments from Section 5, we evaluate a smaller subset of the schedulers available in SAGA: FastestNode, HEFT, CPoP, MaxMin, MinMin, and WBA. Performing these experiments for the rest of the schedulers remains a task for future work. For generating a benchmarking dataset, we use the WfCommons Synthetic Workflow Generator (Coleman et al., 2023) to generate random in-family task graphs and create random networks using a best-fit distribution computed from the real execution trace data. We also use this method to generate initial problem instances for PISA. Then, we adapt PISA’s Perturb implementation as follows:
@@
  • Change Network Node Weight: Same as described in Section 5, except the weight is scaled between the minimum and maximum node speeds observed in the real execution trace data.
@@
  • Change Network Edge Weight: This option is removed since network edge weights are homogeneous and fixed to ensure the average CCR for the given application is a specific value.
@@
  • Change Task Weight: Same as described in Section 5, except the weight is scaled between the minimum and maximum task runtime observed in the real execution trace data.
@@
  • Change Dependency Weight: Same as described in Section 5, except the weight is scaled between the minimum and maximum task I/O size observed in the real execution trace data.
@@
  • Add/Remove Dependency: This option is removed so that the task graph structure remains representative of the real application.
@@
These adjustments to the Perturb implementation allow us to explore application-specific problem instances more realistically.
6.2. Results
@@
In this section we present and discuss some results on two of the workflows analyzed: srasearch and blast. Due to space constraints, full results can be found in Appendix A.3. Srasearch (workflow structure depicted in Figure 5(a)) is a toolkit for interacting with data in the INSDC Sequence Read Archives and blast (workflow structure depicted in Figure 5(b)) is a toolkit for finding regions of similarity between biological sequences.
@@
Figure 5. Example workflow structures for blast and srasearch scientific workflows.
@@
Observe in Figure 5 that while the number of tasks may vary, the structure of both workflows is very rigid. Our new Perturb implementation, though, guarantees that the search space contains only task graphs with appropriate structure. Figures 6 and 7 show the benchmarking and PISA results for srasearch and blast, respectively. First, observe that the benchmarking approach suggests all algorithms (except FastestNode) perform very well on the srasearch applications with a CCR of $1/5$. Using PISA, however, we are able to identify problem instances where WBA performs thousands of times worse than FastestNode! Also, we’re able to find instances where MinMin performs almost twice as bad CPoP. Even among the “good” algorithms, though, we see interesting behavior. Observe the results of HEFT and MaxMin. We are able to find both a problem instance where HEFT performs approximately $20\%$ worse that MaxMin and an instance where MaxMin performs approximately $11\%$ worse than HEFT.
@@
Figure 6. Benchmarking and PISA results for srasearch workflows with different average CCRs (results for CCR of $5$ can be found in Appendix A.3). For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
@@
The results for srasearch seem to imply that CPoP, which appears to perform consistently well for all CCRs tested (the worst case found has a makespan ratio of 1.15) would be the scheduling algorithm of choice for a WFMS designer. Workflow Management Systems do not support just one type of scientific workflow, though, and CPoP’s effectiveness for srasearch workflows may not extend to others. This is true for blast (see results in Figure 7) where CPoP performs generally poorly for all CCRs tested and in one case yields a schedule with 1000 times the makespan of the schedule WBA produces!
@@
Figure 7. Benchmarking and PISA results for blast workflows with different average CCRs (results for CCR of $5$ can be found in Appendix A.3). For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
@@
These results are evidence that the traditional benchmarking approach to comparing algorithms is insufficient even for highly regular, application-specific problem instances. They also have implications for the design of Workflow Management Systems (and other task scheduling systems in IoT, edge computing environments, etc.). It may be reasonable for a WFMS to run a set of scheduling algorithms that best covers the different types of client scientific workflows and computing systems. For example, a WFMS designer might run PISA and choose the three algorithms with the combined minimum maximum makespan ratio. Different methodologies for constructing and comparing such hybrid algorithms is an interesting topic for future work.
7. Conclusion
In this paper, we presented SAGA, a Python framework for running, evaluating, and comparing task scheduling algorithms. We evaluated 15 of the scheduling algorithms implemented in SAGA on 16 datasets and demonstrated how our proposed adversarial analysis method, PISA, provides useful information that traditional benchmarking does not. We also explored how PISA can be used for application-specific scenarios where system designers have some idea of what the target task graphs and compute networks look like ahead of time. We show that, even for this restricted case, PISA is successful in identifying performance boundaries between task scheduling algorithms that the traditional benchmarking approach is not.
@@ 7. Conclusion
There are many directions for future work. First, we plan to extend SAGA to include more algorithms and datasets. Another logical next step is to extend SAGA and the adversarial analysis method to support other problem variants. Due to our particular interest in task scheduling for dynamic environments, we plan to add support for stochastic problem instances (with stochastic task costs, data sizes, computation speeds, and communication costs). It would also be interesting to explore other meta-heuristics for adversarial analysis (e.g., genetic algorithms) and other performance metrics (e.g., throughput, energy consumption, cost, etc.). Finally, the application-specific scenario we explored in Section 6 suggests that an exploration into different methodologies for constructing and comparing hybrid scheduling algorithms using PISA might be fruitful.
Acknowledgements
This work was supported in part by Army Research Laboratory under Cooperative Agreement W911NF-17-2-0196.
@@ Appendix A Appendix
A.1. Scheduling Algorithm Descriptions
To orient the reader, we provide a brief description of each of the algorithms implemented in SAGA and presented in Section 3.
CPoP (Critical Path on Processor) was proposed in the same paper as HEFT (Heterogeneous Earliest Finish Time). Both are list scheduling algorithms with scheduling complexity $O\left(|T|^{2}|V|\right)$. No formal bounds for HEFT and CPoP are known. HEFT works by first prioritizing tasks based on their upward rank, which is essentially the length of the longest chain of tasks (considering average execution and communication times). Then, it greedily schedules tasks in this order to the node that minimizes the task completion time given previously scheduled tasks. CPoP is similar but uses a slightly different priority metric. The biggest difference between the two is that CPoP always schedules critical path tasks (those on the longest chain in the task graph) to execute on the fastest node. Both algorithms were evaluated on random graphs (the process for generating graphs is described in the paper) and on real task graphs for Gaussian Elimination and FFT applications. They were compared against different scheduling algorithms: Mapping Heuristic (similar to HEFT without insertion) (El-Rewini and Lewis, 1990), Dynamic-Level Scheduling (Sih and Lee, 1993a), and Levelized Min Time [^3] We could not find the original paper that proposes this algorithm.. Schedule Length Ratios (makespan scaled by the sum of minimum computation costs of all tasks), speedup, and schedule generation times are reported.
The MinMin, MaxMin, and Duplex list scheduling algorithms may have been proposed many times independently, but exceptionally clear definitions can be found in a paper that compares them to OLB, MET, MCT, and a few meta-heuristic algorithms (e.g., genetic algorithms and simulated annealing) (Braun et al., 2001). MinMin schedules tasks by iteratively selecting the task with the smallest minimum completion time (given previously scheduled tasks) and assigning it to the corresponding node. MaxMin, on the other hand, schedules tasks by iteratively selecting the task with the largest minimum completion time and assigning it to the the corresponding node. Duplex simply runs both MinMin and MaxMin and returns the schedule with the smallest makespan. In the paper, the authors evaluate these algorithms on independent non-communicating tasks with uniformly random costs and heterogeneous compute nodes with uniformly random speeds. MinMin (and therefore also Duplex) is shown to generate schedules with low makespan compared to the other algorithms while relatively high makespans for MaxMin are reported.
FCP (Fast Critical Path) and FLB (Fast Load Balancing) both have a runtime of $O(|T|\log\left(|V|)+|D|\right)$ and were designed for heterogeneous task graphs and heterogeneous node speeds but homogeneous communication strengths. The algorithms were evaluated on three types of task graphs with different structures based on real applications (LU decomposition, Laplace equation solver, and a stencil algorithm) with varied CCR and uniformly random task costs. Both algorithms were shown to perform well compared to HEFT and ERT (Lee et al., 1988) despite their lower schedule generation times.
GDL (Generalized Dynamic Level), also called DLS (Dynamic Level Scheduling), is a variation on list scheduling where task priorities are updated each time a task is scheduled. Due to this added computation in each iteration, the complexity of DLS is $O(|V|^{3}|T|)$ (a factor $|V|$ greater than HEFT and CPoP). GDL was originally designed for the very general unrelated machines model and was shown to outperform HDLFET (Sih and Lee, 1993b) on randomly generated problem instances (though the method used for generating random task graphs is not well-described) and on four real digital signal processing applications (two sound synthesis algorithms, a telephone channel simulator, and a quadrature amplitude modulation transmitter).
OLB (Opportunistic Load Balancing) has a runtime of just $O(|T|)$ and was designed for independent tasks under the unrelated machines model. Probably useful only as a baseline for understanding the performance of other algorithms, OLB schedules tasks in arbitrary order on the earliest available compute node. Its performance has been shown to be significantly worse than MET, MCT, and LBA (Armstrong et al., 1998).
A.2. Case Study: HEFT vs. CPoP
The HEFT and CPoP algorithms were proposed in the same paper by Topcuoglu et al. in (1999) and have remained two of the most popular task scheduling algorithms in decades since. Both are list-scheduling algorithms that use a priority list to determine the order in which tasks are greedily scheduled. In HEFT, each task is scheduled on the node that minimizes its finish time, given previous scheduling decisions. CPoP is similar, but commits to scheduling all critical-path tasks (those on the longest path in the task graph) on the fastest node [^4] This statement is true for the related machines model. In general, CPoP schedules critical-path tasks to the node that minimizes the sum of execution times of critical-path tasks on that node.. The priority function for these algorithms are slightly different from each other, but are both based on task distance (in terms of average compute and communication time) from the start and/or end of the task graph. In HEFT, the priority of a sink task (one with no dependencies) is its average execution time over all nodes in the network. A non-sink task’s priority is the sum of its average execution time over all nodes in the network plus the maximum communication and compute time of its successors. In other words, it’s the average amount of time it takes to execute the task on any node in the network plus the maximum amount of time it takes (in an ideal scenario) to execute the rest of the task graph. In CPoP, the task priorities are computed based on the distance to the end and from the start of the task graph. This difference is a large part of what causes the two algorithms to perform differently on certain problem instances. Observe in Figure 8 a problem instance identified by PISA where HEFT performs approximately $1.55$ times worse than CPoP.
(a) Task Graph
(b) Network
(c) HEFT Schedule
(d) CPoP Schedule
Figure 8. Problem Instance where HEFT performs $\approx 1.55$ times worse than CPoP.
In this instance, the algorithms differ primarily in whether task $A$ or task $C$ has higher priority. In both algorithms, task $B$ has the highest priority and is thus scheduled first on node $2$ (the fastest node). For CPoP, task $C$ must have the highest priority because it is on the critical path $B\rightarrow C$. For HEFT, though, task $A$ has the higher priority than $C$ because it is further from the end of the task graph. As a result, CPoP schedules $C$ on node $2$, which allows task $A$ to be scheduled and executed in parallel on node $3$ (the second fastest node). HEFT, on the other hand, schedules all tasks on node $2$ and thus does not benefit from parallel execution. CPoP succeeds in this instance because it prioritizes tasks that are on the critical path, keeping high-cost tasks/communications on the same node and allowing low-cost tasks/communications to be executed in parallel on other nodes. HEFT greedily schedules high-cost tasks on fast nodes without taking into as much consideration how doing so might affect the rest of the task graph (especially the tasks on the critical path).
This does not mean that CPoP is better than HEFT, though! Observe in Figure 9 a problem instance identified by PISA where CPoP performs approximately $1.88$ times worse than HEFT.
(a) Task Graph
(b) Network
(c) HEFT Schedule
(d) CPoP Schedule
Figure 9. Problem Instance where CPoP performs $\approx 1.88$ times worse than HEFT.
In this instance, the algorithms priorities are actually the same for all tasks. In fact, they schedule both $A$ and $B$ identically. The problem that CPoP faces for this instance is exactly what allows it to succeed in the previous instance: committing to scheduling all critical-path tasks on the fastest node. In this instance, the critical path is $A\rightarrow C$. Thus, CPoP schedules task $C$ on node $2$ (the fastest node) even though it would have finished much faster (due to the communication cost incurred by its dependency on task $A$) on node $1$. HEFT does not have this problem because it does not commit to scheduling all critical-path tasks on the fastest node.
In conclusion, CPoP succeeds when the task has a critical path with low-cost tasks that “mess up” HEFTs priority function and fails when scheduling a task on the fastest node causes a large communication penalty. It’s easy to see that these bad case-scenarios can be modified to make HEFT and CPoP perform arbitrarily worse than each other (e.g., by increasing the number of parallel source tasks in the first example and by changing the network weight between nodes $2$ and $3$ to 00 in the second example). This example highlights one of the benefits of our proposed approach to comparing task scheduling algorithms: it allows us to identify patterns where the algorithms under-/over-perform and then extrapolate those patterns to other problem instances.
A.3. Application-Specific PISA Results
@@
Here we include the Section 6 benchmarking and PISA results for all evaluated scientific workflows. For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
Figure 10. Results for the blast scientific workflow.
Figure 11. Results for the srasearch scientific workflow.
Figure 12. Results for the bwa scientific workflow.
Figure 13. Results for the epigenomics scientific workflow.
Figure 14. Results for the 1000genome scientific workflow.
Figure 15. Results for the montage scientific workflow.
Figure 16. Results for the seismology scientific workflow.
Figure 17. Results for the soykb scientific workflow.

### v3

@@ VII-A Experimental Setup
One of the largest communities of researchers that depend on efficient task scheduling algorithms is that of scientists that use scientific workflows for simulation, experimentation, and more. Scientific workflows are typically big-data scale task graphs that are scheduled to run on cloud or super computing platforms. These researchers typically have little to no control over how their workflows are scheduled, instead relying on Workflow Management Systems (WFMS) like Pegasus [38], Makeflow [39], and Nextflow [40] (to name a just few examples) to handle the technical execution details. For this reason, it is especially important for WFMS developers/maintainers (who choose which scheduling algorithms their system uses) to understand the performance boundaries between different algorithms for the different types of scientific workflows and computing systems their clients use.
@@
SAGA currently has datasets based on nine real-world scientific workflows (blast, bwa, cycles, epigenomics, 1000genome, montage, seismology, soykb, and srasearch). These applications come from a very wide range of scientific domains — from astronomy (montage builds mosaics of astronomical imagery) to biology (1000genome performs human genome reconstruction) to agriculture (cycles is an agroecosystem model). For each workflow, the runtime of each task, input/output sizes in bytes, and speedup factor (compute speed) for each machine are available from public execution trace information [^4] https://github.com/wfcommons/pegasus-instances, https://github.com/wfcommons/makeflow-instances. The inter-node communication rate, however, is not available. We set communication rates to be homogeneous so that the average CCR, or average data sizecommmunication $\frac{\text{average data size}}{\text{commmunication strength}}$, is $\frac{1}{5},\frac{1}{2},1,2,$ or $5$ (resulting in five experiments for each workflow).
@@
Because these scientific workflows are much larger than the workflows used in the experiments from Section VI, we evaluate a smaller subset of the schedulers available in SAGA: FastestNode, HEFT, CPoP, MaxMin, MinMin, and WBA. Performing these experiments for the rest of the schedulers remains a task for future work. For generating a benchmarking dataset, we use the WfCommons Synthetic Workflow Generator [37] to generate random in-family task graphs and create random networks using a best-fit distribution computed from the real execution trace data. We also use this method to generate initial problem instances for PISA. Then, we adapt PISA’s Perturb implementation as follows:
@@
  • Change Network Node Weight: Same as described in Section VI, except the weight is scaled between to the range of speeds observed in the real execution trace data.
@@
  • Change Network Edge Weight: Removed since network edge weights are homogeneous and fixed to enforce a specific CCR.
@@
  • Change Task Weight: Same as described in Section VI, except the weight is scaled to the range of task runtimes observed in the real execution trace data.
@@
  • Change Dependency Weight: Same as described in Section VI, except the weight is scaled between the minimum and maximum task I/O size observed in the real execution trace data.
@@
  • Add/Remove Dependency: Removed so that the task graph structure is representative of the real application.
@@
These adjustments to allow us to explore application-specific problem instances more realistically.
VII-B Results
@@
In this section we present and discuss some results on two of the workflows analyzed: srasearch and blast. Full results can be found in Appendix A-A. Srasearch (workflow structure depicted in Figure 9(a)) is a toolkit for interacting with data in the INSDC Sequence Read Archives and blast (workflow structure depicted in Figure 9(b)) is a toolkit for finding regions of similarity between biological sequences.
@@
Figure 9: Example workflow structures for blast and srasearch scientific workflows.
@@
Observe in Figure 9 that while the number of tasks may vary, the structure of both workflows is very rigid. Our new Perturb implementation, though, guarantees that the search space contains only task graphs with appropriate structure. Figures 10 and 11 show the benchmarking and PISA results for srasearch and blast, respectively (results for some CCRs are excluded due to space constraints and can be found in the full version of the paper [41]). First, observe that the benchmarking approach suggests all algorithms (except FastestNode) perform very well on the srasearch applications with a CCR of $1/5$. Using PISA, however, we are able to identify problem instances where WBA performs thousands of times worse than FastestNode! Also, we’re able to find instances where MinMin performs almost twice as bad CPoP. Even among the “good” algorithms, though, we see interesting behavior. Observe the results of HEFT and MaxMin. We are able to find both a problem instance where HEFT performs approximately $20\%$ worse that MaxMin and an instance where MaxMin performs approximately $11\%$ worse than HEFT.
@@
Figure 10: Benchmarking and PISA results for srasearch workflows with different average CCRs. For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
@@
The results for srasearch seem to imply that CPoP, which appears to perform consistently well for all CCRs tested would be the best choice for a WFMS designer. Workflow Management Systems do not support just one type of scientific workflow, though, and CPoP’s effectiveness for srasearch workflows may not extend to others. This is, in fact, true for blast (see results in Figure 11) where CPoP performs generally poorly for all CCRs tested and in one case yields a schedule with 1000 times the makespan of the schedule WBA produces!
@@
Figure 11: Benchmarking and PISA results for blast workflows with different average CCRs. For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
@@
These results are evidence that the traditional benchmarking approach to comparing algorithms is insufficient even for highly regular, application-specific problem instances. They also have implications for the design of Workflow Management Systems (and other task scheduling systems in IoT, edge computing environments, etc.). It may be reasonable for a WFMS to run a set of scheduling algorithms that best covers the different types of client scientific workflows and computing systems. For example, a WFMS designer might run PISA and choose the three algorithms with the combined minimum maximum makespan ratio. Exploring different methods for constructing and comparing such hybrid algorithms is an interesting topic for future work.
VIII Conclusion
In this paper, we presented SAGA, a Python framework for running, evaluating, and comparing task scheduling algorithms. We evaluated 15 of the scheduling algorithms implemented in SAGA on 16 datasets and demonstrated how our proposed adversarial analysis method, PISA, provides useful information that traditional benchmarking does not. We showed that many algorithms that appear to perform well on benchmarking datasets can perform poorly on adversarially chosen problem instances, effectively presenting new lower bounds on the worst-case performance of algorithms compared to popular baselines. We showed that PISA’s effectiveness extends even to highly restricted, application-specific scenarios. We observed that there exist problem instances where HEFT, a popular task scheduling algorithm, can perform significantly worse than FastestNode, a very simple, generally quite poor scheduling algorithm. Applying this kind of analysis to other pairs of algorithms and exploring how it might be automated is an important and interesting direction for future work.
There are many other directions for future work as well. First, we plan to extend SAGA to include more algorithms and datasets. Another logical next step is to extend SAGA and PISA to support other problem variants (e.g., throughput maximization, energy minimization, etc.). We also plan to develop a framework for publishing the problem instances identified by PISA so that other researchers can use them to evaluate their own algorithms.
@@ VIII Conclusion
Due to our particular interest in task scheduling for dynamic environments, we plan to add support for stochastic problem instances (with stochastic task costs, data sizes, computation speeds, and communication costs). It would also be interesting to explore other meta-heuristics for adversarial analysis (e.g., genetic algorithms) and other performance metrics (e.g., throughput, energy consumption, cost, etc.). Some other directions that could be interesting to explore include ensemble methods (e.g., running multiple algorithms and choosing the best schedule) and online scheduling (e.g., scheduling tasks as they arrive). Finally, the application-specific scenario we explored in Section VII suggests that an exploration into different methodologies for constructing and comparing hybrid scheduling algorithms using PISA might be fruitful.
@@ Appendix A Appendix
A-A Application-Specific PISA Results
@@
Here we include the Section VII benchmarking and PISA results for all evaluated scientific workflows. For each CCR, the top row shows benchmarking results, where gradients indicate makespan ratios on different problem instances in the dataset. All other cells indicate the highest makespan ratio yielding problem instance discovered by PISA.
Figure 12: Results for the blast scientific workflow.
Figure 13: Results for the srasearch scientific workflow.
Figure 14: Results for the bwa scientific workflow.
Figure 15: Results for the epigenomics scientific workflow.
Figure 16: Results for the 1000genome scientific workflow.
Figure 17: Results for the montage scientific workflow.
Figure 18: Results for the seismology scientific workflow.
Figure 19: Results for the soykb scientific workflow.